[Brahms-l] Shift report 20040118 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 18 2004 - 15:46:32 EST
------------------------------------------------------------
Date: Sun Jan 18 15:46:32 2004, Shift: 08:00-16:00
Supervisor: Roy WADA, with: Ionut ARSENE
Log book#15 pages 36 - 36

Shift Summary:
------------------------------------------------------------
1. Few power supply problems. Beam for about 2.5 hours.

2. 9740 - Pedestals
   9741 - TPC pedestals
Data:
   9742, 9743, 9744, 9745(stil running).
  Settings: FS@12deg with D1@A861; MRS@90deg with D5@B2255
  No. of trig3 : >150k
  No. of trig6 : >240k

3. T3 is very close to the trip limit. Anode currents are very high.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 18 15:46:47 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 18 2004 - 15:47:05 EST